### v1

@@ Appendix A Additional proofs
In this section, we provide the proof Proposition 3.6. We begin with Lemma A.1 as a stepping stone. In essence, this results show that the basis computed in Algorithm 1 do not need to be orthonormal, and can always be made an orthonormalization process, and that diagonal blocks can be added while preserving the telescoping factorization. Such comments have been made before, see e.g. [21, Remark 3.3] and [50, Section 4.2].
@@
$\bm{U}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{U}_{1}^{(\ell)},\ldots,\bm{U}_{2^{\ell}}^{(\ell)}\right),\quad\bm{U}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$,
$\bm{V}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{V}_{1}^{(\ell)},\ldots,\bm{V}_{2^{\ell}}^{(\ell)}\right),\quad\bm{V}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$,
@@
$\widetilde{\bm{R}}$ $=\operatorname{blockdiag}\left(\begin{bmatrix}\widetilde{\bm{R}}_{1}&\\ &\widetilde{\bm{R}}_{2}\end{bmatrix},\ldots,\begin{bmatrix}\widetilde{\bm{R}}_{2^{\ell}-1}&\\ &\widetilde{\bm{R}}_{2^{\ell}}\end{bmatrix}\right),$
$\widetilde{\bm{L}}$ $=\operatorname{blockdiag}\left(\begin{bmatrix}\widetilde{\bm{L}}_{1}&\\ &\widetilde{\bm{L}}_{2}\end{bmatrix},\ldots,\begin{bmatrix}\widetilde{\bm{L}}_{2^{\ell}-1}&\\ &\widetilde{\bm{L}}_{2^{\ell}}\end{bmatrix}\right),$
@@
$\bm{U}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{U}_{1}^{(\ell)},\ldots,\bm{U}_{2^{\ell}}^{(\ell)}\right)\quad\text{where }\bm{U}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$,
$\bm{V}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{V}_{1}^{(\ell)},\ldots,\bm{V}_{2^{\ell}}^{(\ell)}\right)\quad\text{where }\bm{V}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$,
@@ Appendix C Bounds for uniform block low-rank approximation
In this section, we consider the class of uniform block low-rank, or $\operatorname{BLR^{2}}$, matrices [4, 59]. In particular, we will provide an error bound for [59, Algorithm 1]. Uniform block low-rank matrices matrices are slightly more general than the class $\operatorname{SSS}(\ell,k)$ defined in Definition 2.5; a uniform BLR matrix allows the sparse remainder matrix $\bm{D}$ to have a different sparsity pattern than diagonal. In particular, $\bm{D}$ is block-diagonal in Definition 2.5, whereas, if one uses a strongly admissible partition for a uniform BLR matrix, $\bm{D}$ is block-tridiagonal. Formally, the class of uniform block low-rank matrices are defined as follows.
@@
$\bm{U}$ $=\operatorname{blockdiag}\left(\bm{U}_{1},\ldots,\bm{U}_{b}\right)\quad\text{where }\bm{U}_{i}\in\mathbb{R}^{m\times k}\text{ has orthonormal columns,}$,
[…]
@@ C.1 Notation
$R_{i}$ =$=\{j\in\{1,\ldots,b\}:(i,j)\in S\},\quad R_{i}^{\prime}=\{j\in\{1,\ldots,b\}:(i,j)\not\in S\}$
$C_{j}$ =$=\{i\in\{1,\ldots,b\}:(i,j)\in S\},\quad C_{j}^{\prime}=\{i\in\{1,\ldots,b\}:(i,j)\not\in S\}.$
@@ C.1.1 Preliminary results
$\operatorname{\mathbb{E}}\Bigl{[}\|\bm{A}-(\bm{U}\bm{X}\bm{V}^{\mathsf{T}}+\bm{D})\|_{\mathsf{F}}^{2}\Bigr{]}\leq(\Gamma_{\textup{r}}+\Gamma_{\textup{c}})(1+\Gamma_{\textup{d}})\cdot\inf\limits_{\bm{B}\in\operatorname{BLR^{2}}(b,m,k,S)}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}.$
@@ C.1.3 Computing the block-sparse matrix
$\bm{U}_{1},\ldots,\bm{U}_{b}\quad$ where $\text{where }\bm{U}_{i}\in\mathbb{R}^{m\times k}\text{ has orthonormal columns,}$,
$\bm{V}_{1},\ldots,\bm{V}_{b}\quad$ where $\text{where }\bm{V}_{i}\in\mathbb{R}^{m\times k}\text{ has orthonormal columns}.$
@@ C.1.4 The algorithm
12:   $\bm{D}_{i,j}=(\bm{I}-\bm{U}_{i}(\bm{U}_{i})^{\mathsf{T}})\widetilde{\bm{Y}}_{i}(\widetilde{\bm{\Omega}}_{R_{i}}^{\dagger}+\bm{U}_{i}(\bm{U}_{i})^{\mathsf{T}}[(\bm{I}-\bm{V}_{j}(\bm{V}_{j})^{\mathsf{T}})\widetilde{\bm{Z}}_{j}(\widetilde{\bm{\Psi}}_{C_{j}})^{\dagger}]^{\mathsf{T}}$
@@
$\mathbb{E}\Bigl{[}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\Bigr{]}\leq\Gamma\cdot\inf\limits_{\bm{C}\in\operatorname{BLR^{2}}(b,m,k,S)}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2},\quad\text{where}\quad\Gamma=O(1).$

### v2

@@ Appendix A Additional proofs
In this section, we provide the proof Proposition 3.6. We begin with Lemma A.1 as a stepping stone. In essence, this results show that the basis computed in Algorithm 1 do not need to be orthonormal, and can always be made an orthonormalization process, and that diagonal blocks can be added while preserving the telescoping factorization. Such comments have been made before, see e.g. [23, Remark 3.3] and [52, Section 4.2].
@@
$\bm{U}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{U}_{1}^{(\ell)},\ldots,\bm{U}_{2^{\ell}}^{(\ell)}\right),\quad\bm{U}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$
$\bm{V}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{V}_{1}^{(\ell)},\ldots,\bm{V}_{2^{\ell}}^{(\ell)}\right),\quad\bm{V}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$
@@
[…]
@@
$\bm{U}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{U}_{1}^{(\ell)},\ldots,\bm{U}_{2^{\ell}}^{(\ell)}\right)\quad\text{where }\bm{U}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$
$\bm{V}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{V}_{1}^{(\ell)},\ldots,\bm{V}_{2^{\ell}}^{(\ell)}\right)\quad\text{where }\bm{V}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$
@@ Appendix C Bounds for uniform block low-rank approximation
In this section, we consider the class of uniform block low-rank, or $\operatorname{BLR^{2}}$, matrices [4, 61]. In particular, we will provide an error bound for [61, Algorithm 1]. Uniform block low-rank matrices matrices are slightly more general than the class $\operatorname{SSS}(\ell,k)$ defined in Definition 2.5; a uniform BLR matrix allows the sparse remainder matrix $\bm{D}$ to have a different sparsity pattern than diagonal. In particular, $\bm{D}$ is block-diagonal in Definition 2.5, whereas, if one uses a strongly admissible partition for a uniform BLR matrix, $\bm{D}$ is block-tridiagonal. Formally, the class of uniform block low-rank matrices are defined as follows.
@@
$\bm{U}$ $=\operatorname{blockdiag}\left(\bm{U}_{1},\ldots,\bm{U}_{b}\right)\quad\text{where }\bm{U}_{i}\in\mathbb{R}^{m\times k}\text{ has orthonormal columns,}$
$\bm{V}$ $=\operatorname{blockdiag}\left(\bm{V}_{1},\ldots,\bm{V}_{b}\right)\quad\text{where }\bm{V}_{i}\in\mathbb{R}^{m\times k}\text{ has orthonormal columns, and}$
@@ C.1 Notation
$R_{i}$ $=\{j\in\{1,\ldots,b\}:(i,j)\in S\},\quad R_{i}^{\prime}=\{j\in\{1,\ldots,b\}:(i,j)\not\in S\}$
$C_{j}$ $=\{i\in\{1,\ldots,b\}:(i,j)\in S\},\quad C_{j}^{\prime}=\{i\in\{1,\ldots,b\}:(i,j)\not\in S\}.$
@@ C.1.1 Preliminary results
$\operatorname{\mathbb{E}}\Bigl{[}\|\bm{A}-(\bm{U}\bm{X}\bm{V}^{\mathsf{T}}+\bm{D})\|_{\mathsf{F}}^{2}\Bigr{]}\leq(\Gamma_{\textup{r}}+\Gamma_{\textup{c}})(1+\Gamma_{\textup{d}})\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{B}\in\operatorname{BLR^{2}}(b,m,k,S)}}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}.$
@@ C.1.3 Computing the block-sparse matrix
$\bm{U}_{1},\ldots,\bm{U}_{b}\quad$ $\text{where }\bm{U}_{i}\in\mathbb{R}^{m\times k}\text{ has orthonormal columns,}$
$\bm{V}_{1},\ldots,\bm{V}_{b}\quad$ $\text{where }\bm{V}_{i}\in\mathbb{R}^{m\times k}\text{ has orthonormal columns}.$
@@ C.1.4 The algorithm
[…]
@@
$\mathbb{E}\Bigl{[}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\Bigr{]}\leq\Gamma\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{BLR^{2}}(b,m,k,S)}}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2},\quad\text{where}\quad\Gamma=O(1).$
Appendix D The existence of an optimal HSS approximation
Unlike HODLR matrices, where the different levels of the hierarchy are independent, HSS matrices impose nestedness conditions across different levels. The nestedness conditions make it less clear why an optimal approximation from the set $\operatorname{HSS}(L,k)$ should exist. We are not aware of any previous results that guarantee the existence of an optimal HSS approximation. Therefore, for completeness, we include a proof here.
Theorem D.1.
For any $\bm{A}\in\mathbb{R}^{2^{L+1}k\times 2^{L+1}k}$, there exists $\widetilde{\bm{B}}\in\operatorname{HSS}(L,k)$ so that
$\|\bm{A}-\widetilde{\bm{B}}\|_{\mathsf{F}}=\inf\limits_{\bm{B}\in\operatorname{HSS}(L,k)}\|\bm{A}-\bm{B}\|_{\mathsf{F}}.$
Proof D.2.
The function $g(\bm{B}):=\|\bm{A}-\bm{B}\|_{\mathsf{F}}$ is a continuous and coercive function. To ensure the existence of a minimum it is therefore sufficient to prove that the set $\operatorname{HSS}(L,k)$ is closed [13, Proposition A.8].
Any $\bm{B}\in\operatorname{HSS}(L,k)$ admits a telescoping factorization with telescoping rank $k$. By [23, Section 3.3] any matrix that admits a telescoping factorization with telescoping rank at most $k$ is also an HSS matrix with HSS rank at most $k$, and by [57, Section 2.2] a matrix is HSS if and only if all HSS block columns and rows, at every level, have rank at most $k$.
Enumerate the HSS block columns and rows from $1$ to $S$, where $S$ is the total number of HSS block columns and rows. Define the function $f:\mathbb{R}^{N\times N}\to\mathbb{R}^{S}_{\geq 0}$ so that $f(\bm{B})_{i}$ equals the $(k+1)^{\text{th}}$ singular value of the $i^{\text{th}}$ HSS block column/row. Note that by the stability of the singular values [14, Problem 111.6.13], $f$ is a vector-valued function whose components are all continuous, so $f$ is also continuous.
By the characterization above we have
$\operatorname{HSS}(L,k)=\{\bm{B}\in\mathbb{R}^{2^{L+1}k\times 2^{L+1}k}:f(\bm{A})=\bm{0}\},$
which is a closed set by the continuity of $f$.